OpenStudy (anonymous):

Use the formula for S_n to find the sum of the first five terms of the geometric sequence 12,-4,4/3,-4/9,...

OpenStudy (anonymous):

r = -4/12 = -1/3 a = 12 Sn = an - a ------ r - 1 tn = ar^n-1 We will first find tn, i.e. t5; the fifth term of the sequence. t5 = 12(-1/3)^5-1 t5 = 12(-1/3)^4 t5 = 12(0.012345679) t5 = 0.148 Now we can use the geometric series formula to find the sum of the first 5 terms. Sn = 0.148 - 12 --------- (-1/3)-1 Sn = -11.852 -------- -1.33 Sn = 8.911 Therefore, the sum of the first 5 terms of the series is 8.911.

OpenStudy (anonymous):

so u can see the pattern is dividing by negative three each time 12, -4, 4/3, -4/9, 4/27 now we have five terms just plus them altogether. It turns out to be 9.037 and the 037 repeats.

OpenStudy (anonymous):

That was incorrect. It was 244/27.

OpenStudy (anonymous):

yeah so 9.037

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

what about 03,-6,6/5,-6/5,..? I keep trying and get the wrong answer. dividing by -5 right?

OpenStudy (anonymous):

wait the numbers are 3, -6, 6/5, -6/5 right?

OpenStudy (anonymous):

-6/25 is the last one.

OpenStudy (anonymous):

but it's not dividing by -5 because the first number is 3 and if 3 is divided by -5 it is -3/5 not -6 so it's not -5.

OpenStudy (anonymous):

first number 30. sorry didnt see that mistake.

OpenStudy (anonymous):

so it's the first five numbers will be 30, -6, 6/5, -6/25, 6/125 and so we just add them altogether to get 25.008.

OpenStudy (anonymous):

thnx that makes sense:)

OpenStudy (anonymous):

np :)